# The Edge-of-Reach Problem in Offline Model-Based Reinforcement Learning

### v1

@@ A.3 Comparison Between Model-Free and Model-Based Methods
Table 5: A summary of the comparison between model-free and model-based offline RL in relation to the conditions on $Q$-learning.
@@
Model-free Possible, common in practice Possible, but uncommon in practice $\rightarrow$ Action condition violation
Model-based Not possible Possible, common in practice $\rightarrow$ State condition violation
In offline model-free RL, $\mathcal{D}=\mathcal{D}_{\textrm{offline}}$, with $\pi^{\mathcal{D}}=\pi^{\beta}$. For the settings we consider, $\mathcal{D}_{\textrm{offline}}$ consists of full trajectories and therefore will not violate the state condition. However, this may happen in a more general setting with $\mathcal{D}_{\textrm{offline}}$ containing truncated trajectories. By contrast, the mismatch between $\pi$ (used to sample $a^{\prime}$ in $Q$-learning) and $\pi^{\beta}$ (used to sample $a$ in the dataset $\mathcal{D}_{\textrm{offline}}$) often does lead to significant violation of the action condition. This exacerbates the overestimation bias in $Q$-learning (see Section 3), and can result in pathological training dynamics and $Q$-value explosion over training (Kumar et al., 2019).
Appendix B Additional Figures
B.1 Comparison to Prior Approaches
In Figure 6, we plot the dynamics uncertainty-based penalty used in MOPO (Yu et al., 2020) against the effective penalty in RAVL of variance of the value ensemble. The positive correlation between MOPO’s penalty and the edge-of-reach states targeting penalty of RAVL helps to explain why prior methods may work despite not considering the crucial edge-of-reach problem.
Figure 6: We find that the dynamics uncertainty-based penalty used in MOPO (Yu et al., 2020) is positively correlated with the variance of the value ensemble of RAVL, suggesting prior methods may unintentionally address the edge-of-reach problem. Pearson correlation coefficients are 0.49, 0.43, and 0.27 for Hopper-mixed, Walker2d-medexp, and Halfcheetah-medium respectively.
B.2 Visualizations of Rollouts Through Training
In Figure 7, we provide a visualization of the rollouts sampled over training in the simple environment for each of the algorithms analyzed in Figure 4 (see Section 5). This accompanies the discussion of the behavior of SAC during training in Section 5.2.
Figure 7: A visualization of the rollouts sampled over training on the simple environment in Section 5. We note the pathological behavior of the baseline SAC, and the success of the ideal intervention SAC-OraclePatch, and our practically realizable method RAVL.
@@ Appendix C Implementation Details
In this section, we give full implementation details for RAVL.
@@ C.1 Algorithm
As is typical in offline model-based reinforcement learning (Yu et al., 2020; Lu et al., 2022; Kidambi et al., 2020; Sun et al., 2023), we use short MBPO (Janner et al., 2019) style rollouts. The agent is trained using the EDAC (An et al., 2021) losses, a conservative ensemble-based algorithm based on SAC (Haarnoja et al., 2018). Our transition model follows the standard setup in model-based offline algorithms, being realized as a deep ensemble (Chua et al., 2018) and trained via maximum likelihood estimation. This is summarized in the pseudocode for offline model-based methods given in Algorithm 1.
@@ C.2 Hyperparameters
For the D4RL (Fu et al., 2020) MuJoCo datasets presented in Section 7.2, we sweep over the following hyperparameters and list the optimal choices in Table 6.
@@
  • (MBPO) Model rollout length $k$, in the range $\{1,5\}$
  • (MBPO) Real-to-synthetic data ratio $r$, in the range $\{0.05,0.5\}$
The remaining model-based and agent hyperparameters are given in Table 7. Almost all environments use a small $N_{\textrm{critic}}=10$ similar to An et al. (2021), with only the Hopper-medexp dataset needing $N_{\textrm{critic}}=30$. Our implementation is based on the Clean Offline Reinforcement Learning (CORL, Tarasov et al. (2022)) repository, released at https://github.com/tinkoff-ai/CORL under an Apache-2.0 license. Our algorithm takes approximately 1 day to run on the standard D4RL datasets using an A100 GPU for the full number of epochs.
Table 6: Variable hyperparameters for RAVL used in D4RL MuJoCo locomotion tasks.
@@
HalfCheetah medium 10 1 5 0.05
mixed 10 100 5 0.05
medexp 10 1 5 0.5
Hopper medium 10 100 1 0.5
mixed 10 10 1 0.5
medexp 50 100 1 0.5
Walker2d medium 10 10 1 0.5
mixed 10 10 1 0.5
medexp 10 1 1 0.5
@@ C.3 Pixel-Based Hyperparameters
For the V-D4RL (Lu et al., 2023) DeepMind Control Suite datasets presented in Section 7.3, we use the default hyperparameters for the Offline DV2 algorithm, which are given in Table 8. We found keeping the uncertainty weight $\lambda=10$ improved performance over $\lambda=0$ which shows RAVL can be combined with dynamics penalty-based methods.
@@
We swept over $\{2,5,20\}$ for $\textrm{N}_{\textrm{critic}}$ but found a single value of $5$ worked well for all environments we consider.
Appendix D Proof of Error Propagation Result
In this section, we provide a proof of Proposition 1. Our proof follows analogous logic to the error propagation result of Kumar et al. (2019).
Proof.
$\zeta_{j}(s_{t},a_{t})$ $=|Q_{j}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (4)
$=\delta_{j}(s_{t},a_{t})+\gamma\zeta_{j-1}(s_{t+1},a_{t+1})$ (7)
$=\delta_{j}(s_{t},a_{t})+\gamma\epsilon$ (8)
Appendix E Summary of Setups Used in Comparisons
@@
Throughout the paper, we compare several different setups in order to identify the true underlying issues in model-based offline RL. We provide a summary of them in Table 9. More comprehensive descriptions of each are given in the main text and in relevant table and figure captions.
@@
Table 9: We summarize the various setups used for comparisons throughout the paper. ‘*’ denotes application to the simple environment (see Section 5). All methods use $k$-step rollouts from the offline dataset (or from a fixed starting state distribution in the case of the simple environment). Use of the MOPO penalty refers to implementing the method in Yu et al. (2020).
Dynamics Model Agent Figures and Tables
@@
Learned Ensemble ✓(MOPO) $2$ 00 Table 1, Figure 6
Oracle True n/a $2$ 00 Table 1
MBPO (Offline) Ensemble $2$ 00 Tables 2, Figure 3
RAVL (Ours) Ensemble $>2$ $>0$ Tables 2, 3, Figure 3
SAC (Truncated Horizon)* True n/a $2$ 00 Figures 4, 7
RAVL (Ours)* True n/a $>2$ $>0$ Figures 4, 5, 7

### v2

@@ A.3 Comparison Between Model-Free and Model-Based Methods
In offline model-free RL, $\mathcal{D}=\mathcal{D}_{\textrm{offline}}$, with $\pi^{\mathcal{D}}=\pi^{\beta}$. For the settings we consider, $\mathcal{D}_{\textrm{offline}}$ consists of full trajectories and therefore will not violate the state condition. However, this may happen in a more general setting with $\mathcal{D}_{\textrm{offline}}$ containing truncated trajectories. By contrast, the mismatch between $\pi$ (used to sample $a^{\prime}$ in $Q$-learning) and $\pi^{\beta}$ (used to sample $a$ in the dataset $\mathcal{D}_{\textrm{offline}}$) often does lead to significant violation of the action condition. This exacerbates the overestimation bias in $Q$-learning (see Section 7), and can result in pathological training dynamics and $Q$-value explosion over training [18].
This comparison between model-based and model-free is summarized in LABEL:tab:conditions1
Table 3: A summary of the comparison between model-free and model-based offline RL in relation to the conditions on $Q$-learning as described in Appendix A.
@@
Model-free Yes, common in practice Yes, but uncommon in practice $\rightarrow$ Action condition violation
Model-based No Yes, common in practice $\rightarrow$ State condition violation
Appendix B Error Propagation Result
Proposition 1 (Error propagation from edge-of-reach states). Consider a rollout of length $k$, $(s_{0},a_{0},s_{1},\dots,s_{k})$. Suppose that the state $s_{k}$ is edge-of-reach and the approximate value function $Q^{j}(s_{k},\pi(s_{k}))$ has error $\epsilon$. Then, standard value iteration will compound error $\gamma^{k-t}\epsilon$ to the estimates of $Q^{j+1}(s_{t},a_{t})$ for $t=1,\dots,k-1$. (Proof in Appendix B.)
B.1 Proof
In this section, we provide a proof of Proposition 1. Our proof follows analogous logic to the error propagation result of Kumar et al. [18].
Proof.
$\zeta_{j}(s_{t},a_{t})$ $=|Q_{j}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|$ (4)
$=\delta_{j}(s_{t},a_{t})+\gamma\zeta_{j-1}(s_{t+1},a_{t+1})$ (7)
$=\delta_{j}(s_{t},a_{t})+\gamma\epsilon$ (8)
B.2 Extension to Stochastic Environments and Practical Implementation
With stochastic transition models (e.g. Gaussian models), we may have the case that no state will truly have zero density, in which case we relax the definition of edge-of-reach states slightly (see Section 3.3) from $\rho_{t,\pi}(s)=0$ to $\rho_{t,\pi}(s)<\epsilon$ for some small $\epsilon$.
During optimization, in practice, model rollouts are sampled in minibatches and thus the above error propagation effect will occur on average throughout training. An analogous model-free statement has been given in Kumar et al. [18]; however, its significance in the context of model-based methods was previously not considered.
@@ Appendix C Implementation Details
In this section, we provide full implementation details for RAVL.
@@ C.1 Algorithm
We use the base model-based procedure as given in Algorithm 1 and shared across model-based offline RL methods [36, 24, 14, 29]. This involves using short MBPO-style [13] model rollouts to train an agent based on SAC [11]. We modify the SAC agent with the value pessimism losses of EDAC [1]. Our dynamics model follows the standard setup in model-based offline algorithms, being realized as a deep ensemble [5] and trained via maximum likelihood estimation.
@@ C.2 Hyperparameters
For the D4RL [7] MuJoCo results presented in Table 2, we sweep over the following hyperparameters and list the choices used in Table 4. “Base” refers to the shared base procedure in model-based offline RL shown in Algorithm 1.
@@
  • (Base) Model rollout length $k$, in the range $\{1,5\}$
  • (Base) Real-to-synthetic data ratio $r$, in the range $\{0.05,0.5\}$
The remaining model-based and agent hyperparameters are given in Table 5. Almost all environments use a small $N_{\textrm{critic}}=10$, with only the Hopper-medexp dataset needing $N_{\textrm{critic}}=30$.
For the uncertainty-free dynamics model experiments in Table 1 we take the same hyperparameters as for the learned dynamics model for $N$, $k$, and $r$, and try different settings for $\eta\in\{1,10,100,200\}$. The choices used are shown in Table 4. Note that for existing methods based on dynamics penalties, the analogous hyperparameter (penalty weighting) has no effect under the true dynamics model as the penalties will always be zero (since dynamics uncertainty is zero). For the experiments in Figure 1 with intermediate dynamics model accuracies, we tune the ensemble diversity coefficient $\eta$ over $\eta\in\{1,10,100,200\}$ for RAVL, and analogously we tune the uncertainty penalty coefficient $\lambda$ over $\lambda\in\{1,5,10,100,200\}$ for MOPO.
Our implementation is based on the Clean Offline Reinforcement Learning (CORL, Tarasov et al. [33]) repository, released at https://github.com/tinkoff-ai/CORL under an Apache-2.0 license. Our algorithm takes on average 6 hours to run using a V100 GPU for the full number of epochs.
Table 4: Variable hyperparameters for RAVL used in D4RL MuJoCo locomotion tasks with the learned dynamics (see Table 2). With the uncertainty-free dynamics (see Table 1), we tune only $\eta$, with the settings used shown in brackets.
@@
      HalfCheetah       random       10       10 (1)       5       0.5
      medium       10       1 (1)       5       0.05
      mixed       10       100 (1)       5       0.05
      medexp       10       1 (1)       5       0.5
      Hopper       random       10       100 (1)       5       0.05
      medium       10       100 (10)       1       0.5
      mixed       10       10 (10)       1       0.5
      medexp       30       100 (10)       1       0.5
      Walker2d       random       10       1 (100)       5       0.05
      medium       10       10 (100)       1       0.5
      mixed       10       10 (100)       1       0.5
      medexp       10       1 (200)       1       0.5
@@ C.3 Pixel-Based Hyperparameters
For the V-D4RL [25] DeepMind Control Suite datasets presented in Table 9, we use the default hyperparameters for the Offline DV2 algorithm, which are given in Table 6. We found keeping the uncertainty weight $\lambda=10$ improved performance over $\lambda=0$ which shows RAVL can be combined with dynamics penalty-based methods.
@@
We used a hyperparameter sweep over $\{2,5,20\}$ for $\textrm{N}_{\textrm{critic}}$ but found a single value of $5$ worked well for all environments we consider.
Appendix D Additional Tables
D.1 Existing Methods’ Penalties All Collapse to Zero Under the True Dynamics
For completeness, in Table 7 we include the expressions for the various dynamics penalties proposed by the most popular model-based approaches. These penalties are all based on estimates of dynamics model uncertainty. Since there is no uncertainty under the true model, all these penalties collapse to zero with the true dynamics.
Table 7: We show here how all existing dynamics penalized offline MBRL algorithms reduce to the same base procedure when estimated (epistemic) model uncertainty is zero (as it is with the true dynamics).
Offline MBRL Algorithm Penalty Penalty with the true dynamics
MOPO [36] ${\max}_{i=1,\dots,N}||\Sigma^{i}_{\phi}(s,a)||_{\mathrm{F}}$ 0
MOReL [14] ${\max}_{1\leq i,j\leq N}||\mu^{i}_{\phi}(s,a)-\mu^{j}_{\phi}(s,a)||_{2}$ 0
MOBILE [29] $\textrm{Std}_{i=1,\dots,N}\{\hat{T}^{i}(s,a)\}$ 0
D.2 Model Errors Cannot Explain the $Q$-value Overestimation
In Figure 3 we observe that the base offline model-based procedure results in $Q$-values exploding unboundedly to beyond $10^{10}$ over training. Existing methods attribute this to dynamics model errors and the consequent exploitation of state-action space areas where the model incorrectly predicts high reward.
In Table 8 we examine the rewards collected by an agent in the learned dynamics model. Under the ‘model errors’ explanation, we would expect to see rewards sampled by the agent in the learned model to be significantly higher than under the true environment. Furthermore, to explain the $Q$-values being on the order of $10^{10}$, with $\gamma=0.99$ the agent would need to sample rewards on the order of $10^{8}$. However, the per-step rewards sampled by the agent are on the order of $10^{0}$ (and are not higher than with the true environment).
Model errors and model exploitation therefore cannot explain the explosion in $Q$-values seen over training. By contrast, the edge-of-reach problem exactly predicts this extreme value overestimation behavior.
Table 8: Per-step rewards with the base offline model-based procedure (see Algorithm 1). Rewards in model rollouts are close to those with the true dynamics, showing that model exploitation could not explain the subsequent value overestimation. Further, it indicates (as in Lu et al. [24], Janner et al. [13]) that the model is largely accurate for short rollouts and hence is unlikely to be vulnerable to model exploitation.
            Environment             Model reward             True reward
            Hc-med             4.89$\pm$1.27             5.01$\pm$1.09
            Hop-mixed             2.37$\pm$1.05             2.44$\pm$1.02
            W2d-medexp             3.96$\pm$1.41             3.96$\pm$1.41
Appendix E Additional Visualizations
E.1 Comparison to Prior Approaches
In Figure 6, we plot the dynamics uncertainty-based penalty used in MOPO [36] against the effective penalty in RAVL of variance of the value ensemble. The positive correlation between MOPO’s penalty and the edge-of-reach states targeting penalty of RAVL indicates why prior methods may work despite not considering the crucial edge-of-reach problem.
Figure 6: We find that the dynamics uncertainty-based penalty used in MOPO [36] is positively correlated with the variance of the value ensemble of RAVL, suggesting prior methods may unintentionally address the edge-of-reach problem. Pearson correlation coefficients are 0.49, 0.43, and 0.27 for Hopper-mixed, Walker2d-medexp, and Halfcheetah-medium respectively.
E.2 Behaviour of Rollouts Over Training
In Figure 7, we provide a visualization of the rollouts sampled over training in the simple environment for each of the algorithms analyzed in Figure 4 (see Section 4). This accompanies the discussion in Section 4.2 of the behavior over training.
Figure 7: A visualization of the rollouts sampled over training on the simple environment in Section 4. We note the pathological behavior of the baseline, and then the success of the ideal intervention Base-OraclePatch, and our practically realizable method RAVL.
Appendix F Evaluation on the Pixel-Based V-D4RL Benchmark
The insights from RAVL also lead to improvements in performance on the challenging pixel-based benchmark V-D4RL [25]. The base procedure in this pixel-based setting uses the Offline DreamerV2 algorithm of training on trajectories in a learned latent space.
We observe in Table 9 that RAVL gives a strong boost in performance on the medium and medexp level datasets, while helping less in the more diverse random and mixed datasets. This observation fits with our intuition of the edge-of-reach problem: the medium and medexp level datasets are likely to have less coverage of the state space, and thus we would expect them to suffer more from edge-of-reach issues and the ‘edge-of-reach seeking behavior’ demonstrated in Section 4.2. The performance improvements over Offline DreamerV2 and other model-based baselines including LOMPO [26] therefore suggests that the edge-of-reach problem is general and widespread.
We also note that these results are without the ensemble diversity regularizer from EDAC [1] (used on the D4RL benchmark), which we anticipate may further increase performance.
Table 9: We show that RAVL extends to the challenging pixel-based V-D4RL benchmark, suggesting the out-of-reach problem is also present in the latent-space setting used by the base procedure for pixel-based algorithms. Mean and standard deviation given over 6 seeds.
      V-D4RL Environment       Offline DV2       LOMPO       RAVL (Ours)
      Walker-walk       random       28.7$\pm$13.0       21.9$\pm$8.1       29.7$\pm$2.4
      mixed       56.5$\pm$18.1       34.7$\pm$19.7       49.3$\pm$11.3
      medium       34.1$\pm$19.7       43.4$\pm$11.1       48.6$\pm$8.7
      medexp       43.9$\pm$34.4       39.2$\pm$19.5       47.6$\pm$18.3
      Cheetah-run       random       31.7$\pm$2.7       11.4$\pm$5.1       36.9$\pm$5.0
      mixed       61.6$\pm$1.0       36.3$\pm$13.6       57.6$\pm$1.4
      medium       17.2$\pm$3.5       16.4$\pm$8.3       18.3$\pm$4.3
      medexp       10.4$\pm$3.5       11.9$\pm$1.9       12.6$\pm$4.4
      Overall       35.5$\pm$12.0       26.9$\pm$10.9       37.6$\pm$7.0
Appendix G Runtime and Hyperparameter Sensitivity Ablations
G.1 Runtime
We compare RAVL to SOTA method MOBILE. The additional requirement for RAVL compared to MOBILE is that RAVL requires increasing the $Q$-ensemble size: from $N=2$ in MOBILE to $N=10$ in RAVL (for all environments except one where we use $N=30$). The additional requirement for MOBILE compared to RAVL is that MOBILE requires multiple additional forwards passes through the model for each update step (for computing their dynamics uncertainty penalty).
In Table 10 we show that ensemble size can be scaled with very minimal effect on the runtime. Even increasing the ensemble size to $N=100$ (far beyond the maximum ensemble size used by RAVL) only increases the runtime by at most 1%. Table 7 of MOBILE reports that MOBILE’s requirement of additional forward passes increases runtime by around 14%.666 Note that the times reported for EDAC in Table 7 of MOBILE are not representative of practice, since these are with a non-vectorized $Q$-ensemble (confirmed from their open-source code), which would require just a simple change to  5 lines of code.
Table 10: Timing experiments showing that ensemble size for standard vectorized ensembles can be scaled with extremely minimal effect on runtime. Runtime increases by just 1% with $Q$-ensemble size increased from $N=2$ (as used in the base procedure SAC agent) to $N=100$. RAVL uses $N=10$ (for all environments except one where we use $N=30$), meaning the runtime increase compared to the base model-based procedure is almost negligible. Other methods add more computationally expensive changes to the base procedure, hence making RAVL significantly faster than SOTA (see Section G.1).
     Ensemble size ($N$) 2 10 50 100
Time relative to $N=2$ (baseline) ×$\times 1.00\pm 0.03$$\times 1.01\pm 0.04$$\times 1.01\pm 0.02$
G.2 Hyperparameter Sensitivity Ablations
In Table 11 we include results with different settings of RAVL’s ensmble diversity regularizer hyperparameter $\eta$. We note a minimal change in performance with a sweep across two orders of magnitude. This is an indication that RAVL may be applied to new settings without much tuning.
For RAVL’s second hyperparameter of $Q$-ensemble size $N$, we note that in the main benchmarking of RAVL against other offline methods (see Table 2) we use the same value of $N=10$ across all except one environment. This again indicates that RAVL is robust and should transfer to new settings with minimal hyperparameter tuning.
Table 11: Ablations results over RAVL’s diversity regularizer hyperparameter $\eta$.
         Hyperparameter $\eta$          Halfcheetah medium          Halfcheetah mixed
         1.0          $78.7\pm 2.0$          $61.7\pm 5.6$
         10.0          $74.8\pm 1.5$          $73.8\pm 1.9$
         100.0          $72.1\pm 2.2$          $74.9\pm 2.0$
Appendix H Summary of Setups Used in Comparisons
@@
Throughout the paper, we compare several different setups in order to identify the true underlying issues in model-based offline RL. We provide a summary of them in Table 12. More comprehensive descriptions of each are given in the relevant positions in the main text and table and figure captions.
@@
Table 12: We summarize the various setups used for comparisons throughout the paper. ‘*’ denotes application to the simple environment (see Section 4). All methods use $k$-step rollouts from the offline dataset (or from a fixed starting state distribution in the case of the simple environment).
Figures and Tables Name Used Dynamics Model Agent
@@
Table 1, Figure 6 ‘Approximate’ Ensemble ✓(Best of Table 7) $2$ 00
Table 1 ‘True’ True n/a $2$ 00
Tables 8, Figure 3 ‘Base’ Ensemble $2$ 00
Tables 8, 2, Figure 3 ‘RAVL (Ours)’ Ensemble $>2$ $>0$
Figures 4, 7 ‘Base’* True n/a $2$ 00
Figures 4, 5, 7 ‘RAVL (Ours)’* True n/a $>2$ $>0$